OpenStudy (anonymous):

how many atoms are in 8.70 mol of copper (II) Phosphate

OpenStudy (anonymous):

The formula of copper (II) phosphate is \(\sf Cu_3(PO_4)_2\). As you know, 1 mole of any substance consists of \(\sf 6.022×10^{23}\) molecules. From here, you can calculate number of \(\sf Cu_3(PO_4)_2\) molecules in 8.70 moles. \(8.70\times 6.022 \times10^{23}\approx 5.24\times 10^{24}\) molecules Next, from the formula \(\sf Cu_3(PO_4)_2\), you can easily find the number of atoms: \(3+2\times (1+4)=13\). Therefore, the number of atoms is \(\sf 13\times 5.24\times 10^{24}\approx 6.81\times 10^{25}\) atoms